Circular No. 8193

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 CE_{10} (LINEAR) N. Takato, T. Sekiguchi, and J. Watanabe, National Astronomical Ob-

servatory of Japan, obtained nine CCD images with the 8.2-m Subaru telescope of the apparently asteroidal object 2002 CE_{10} (first reported by the LINEAR team, whose discovery observation is given below; originally announced on MPEC 2002-C83 and MPS 50101) on Aug. 22.4 UT that show a very faint, straight tail ∼ 21′′ long in p.a. 212^{o}; the tail is also present on shorter exposures from Aug. 21.5–21.6, when any coma as bright as the tail must have been < 6′′ in diameter. Recent astrometry, orbital elements (T = 2003 June 22, ω = 126^{o}, Ω = 147^{o}, i = 145^{o}, e = 0.79, P = 30.8 yr), and an ephemeris appear on MPEC 2003-R20.

2002 UT  α_{2000}  δ_{2000}  Mag.
Feb. 6.31522  9^{h}15^{m}20.^{s}96  +28^{o}05′14′′.5  19.4

SATELLITES OF NEPTUNE On Sept. 1, S. S. Sheppard, University of Hawaii, reported to the Mi-

nor Planet Center the discovery by D. C. Jewitt, J. Kleyna, and himself of a possible new satellite (mag R = 26) of Neptune with the 8.2-m Subaru reflector at Mauna Kea on Aug. 29.3 and 30.3 UT. He also reported observations from the same nights of S/2002 N 1 (IAUC 8047, MPEC 2003-A75). Later that day, in response to a request by B. G. Marsden (Minor Planet Center), M. Holman, Harvard-Smithsonian Center for Astrophysics, reported observations by J. Kavelaars and himself with the 4-m Blanco reflector at Cerro Tololo on Aug. 21.1 UT of S/2002 N 1, as well as observations on Aug. 19.0 of another Neptunian-satellite candidate that Holman felt might conceivably be identical with a satellite candidate observed in the same manner on 2001 Aug. 11.1. Computations by Marsden showed the new Sheppard and Holman objects to be same and that identity with the 2001 object was a distinct possibility. Holman was then able to locate this object, now designated S/2003 N 1, on two nights in each of July 2003 and Aug. 2002 (with T. Grav and W. Fraser as participating observers). The complete set of observations and a linked orbit by Marsden (a = 0.33 AU, e = 0.27, i = 124^{o}, H = 10.8) are given on MPEC 2003-R19. The 2003 observations (including also some from June 3 and July 29 with the 6.5-m Clay reflector at Las Campanas) of S/2002 N 1 are given on MPEC 2003-R18, together with Marsden’s improved orbit (a = 0.11 AU, e = 0.26, i = 112^{o}, H = 9.8).

2003 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green